OpenStudy (anonymous):

Please please please! I've been stuck on this question and can't figure it out!! Create two sets of three triangle lengths: one set that does create a right triangle and another that does not. Show and explain why the first set does create a right triangle and why the second set does not.

OpenStudy (anonymous):

np. The set that does give a right triangle satisfies a^2 + b^2 = c^2 and one of the classic sets for that is a=3, b=4, c=5, because 3^2 + 4^2 = 9 +16 = 25 = 5^2

OpenStudy (anonymous):

For the second set, you can keep a=3, b=4. For c, you can take any length that is: 0 < c < 5 or 5 < c < 3+4 which is 7. "c" has to be positive obviously since lengths are positive. It can't be 5 because we used that for the first set. And it has to be less than 7 or else it is too long. I'd use c=6.

OpenStudy (anonymous):

The explanation for why the second set does not create a right triangle is that 3^2 + 4^2 does NOT = 6^2

OpenStudy (anonymous):

So, we're done! Does this all make sense to you now?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (anonymous):

You're welcome! And medals are nice!

OpenStudy (anonymous):

If you like my work, just click "best response" next to my posts.